# references.bib

### v1

@@ 4.2 Results
All qualitative results are reported in Table 2. The lowest error is achieved with the architecture of Lang et al., trained on all considered input channels. The best models achieve RMSE values slightly below $60$ Mg/ha, consistent with previous literature.
@@
Role of input features. In order to study the impact of each data source, we conduct a series of experiments. We vary the input features available to each model and alternative feed: all features, only the canopy height (CH), only the 101010\,10m RGB (B02, B03, B04) and NIR (B08) Sentinel-2 bands (RGBN), all Sentinel-2 bands (S2 (all)), all Sentinel-2 bands plus the LC map (LC ), all Sentinel-2 bands plus the DEM (DEM), all Sentinel-2 bands plus the two ALOS-2 PALSAR-2 bands (P2), and features except for the CH. Per-pixel latitude and longitude are always provided as inputs. Each model is evaluated with every feature combination, and in each run the weights with the highest validation score are applied to the held-out test set to obtain the RMSE values reported in Table 2. Additionally we also show feature importance plots obtained with the GBDT model in the Appendix.
Patch size. To study how varying the patch-size affects model performance, we train and evaluate with two different patch sizes, $15\times 15$ and $25\times 25$ pixels, see Table 2.
@@
Binned residuals analysis. Beyond overall test RMSE we also analyse the residual distribution across various AGB intervals, see Figure 6. Besides our baseline models described in Section 4.1 we also include the ESA CCI AGB map, upsampled to 101010\,10m resolution, in this analysis. That comparison is directly relevant, since ESA CCI also uses GEDI to calibrate their model, according to their Algorithm Theoretical Basis Document [ESACCI, p.23]. We point out that the distribution of AGB values is skewed: most ($50$%) of all GEDI labels fall into the first bin ($[0,50]$). Density scatter plots can be found in the Appendix.
Qualitative results inspection. For a visual impression, Figure 5 shows the predictions of the best-performing model versus the CCI maps for two Sentinel-2 tiles: one in Ghana (30NXM) and one in Paraguay (21KVQ).
Figure 5: Maps predicted by the best model vs. ESA CCI maps, for two Sentinel-2 tiles: 30NXM (top) and 21KVQ (bottom). Small offsets are due to Google Earth using a different map projection.
Table 2: Mean test RMSE ($\downarrow$) and associated standard deviation per model, with various inputs. Crosses denote the presence of a feature, values in brackets denote patch size. RMSE values in the lowest 10% of the range are colour-coded, with darker blue indicating better results.
@@
× 69.88 $\pm$0.17 69.50 $\pm$0.13 69.49 $\pm$0.11 68.99 $\pm$0.15 68.73 $\pm$0.22 68.43 $\pm$0.10 68.04 $\pm$0.09
× 74.74 $\pm$0.23 65.35 $\pm$0.16 65.06 $\pm$0.26 63.25 $\pm$0.31 62.11 $\pm$0.52 61.73 $\pm$0.30 60.39 $\pm$0.22
× 70.93 $\pm$0.35 64.16 $\pm$0.22 63.92 $\pm$0.40 61.87 $\pm$0.14 60.94 $\pm$0.22 60.46 $\pm$0.15 60.68 $\pm$0.46
× × 67.52 $\pm$0.15 62.57 $\pm$0.22 62.48 $\pm$0.28 61.17 $\pm$0.23 60.30 $\pm$0.16 59.72 $\pm$0.28 58.80 $\pm$0.08
× × 68.38 $\pm$0.15 63.30 $\pm$0.24 63.92 $\pm$0.34 61.61 $\pm$0.15 60.59 $\pm$0.10 60.05 $\pm$0.13 59.03 $\pm$0.09
× × 68.69 $\pm$0.27 62.19 $\pm$0.35 61.94 $\pm$0.21 60.40 $\pm$0.14 59.24 $\pm$0.18 59.08 $\pm$0.05 58.04 $\pm$0.13
× × × × 65.63 $\pm$0.08 61.47 $\pm$0.21 61.34 $\pm$0.03 59.85 $\pm$0.12 58.97 $\pm$0.14 58.64 $\pm$0.08 57.87 $\pm$0.11
× × × × × 63.64 $\pm$0.06 61.29 $\pm$0.13 61.22 $\pm$0.27 60.02 $\pm$0.18 59.20 $\pm$0.14 58.67 $\pm$0.08 58.16 $\pm$0.19
Figure 6: Binned residuals over the test set for the best-performing version of each architecture, and for the ESA CCI predictions.
@@ 5 Discussion
Results. The input features study confirms that including more input features is beneficial, and that models with more parameters tend to better capture the complex relationship between input features and AGB. The best result overall is achieved by the architecture of Lang et al., with a patch size of $25\times 25$ and all input channels. Thus, it makes sense to assemble a dataset with a wider range of observations. In particular the results underline that, while canopy height and biomass are strongly correlated, it is not straightforward to accurately derive the latter only from the former.
Figure 5 illustrates how the high spatial resolution of our estimates is able to preserve spatial details that are not captured in the CCI product, meaning that they would not be restored by conventional upscaling of existing maps. Given the limited spatial context of most existing biomass retrieval methods, it is interesting that increasing the context window yields better results. At present it is unclear whether this is due to real context information or simply a data augmentation effect. Consistent with the literature, the residuals analysis highlights the commonly observed under-estimation of higher biomass values, although the comparison to CCI suggests that our deep learning methods mitigate that effect.
Limitations. Despite the nominal 10 m resolution of our estimates, the effective spatial resolution at which biomass is identified is lower: we actually learn to predict, at each pixel, the highest biomass within a $\approx 25$m diameter (the size of the GEDI footprint). Another point to keep in mind is that the spatial distribution of GEDI footprints is not uniform: the sampling gets progressively denser as one approaches towards the borders of the observed region at ±$\pm 51.6^{\circ}$.
@@ 6 Conclusion
We provide a ML-ready, unrestricted, easily accessible dataset for high-resolution biomass estimation from remote sensing data. This dataset consists of $\approx 16\cdot 10^{6}$ patches, covering a diverse and representative selection of geographic locations with a total of $\approx 500\cdot 10^{6}$ hectares and two different years, $2019$ and $2020$. Furthermore, we complement the dataset with a suite of baseline architectures and trained weights for others to build upon. We hope that our dataset may help to reduce the geographical and ecosystem-related biases of biomass models, enhance the accuracy of future biomass maps, and perhaps contribute towards a regular, operational high-resolution monitoring system.
@@ 7 Acknowledgments
Supplementary Material
@@ 7.1 Hosting, licensing, and maintenance plan
The raw data, dense predictions, and model weights are hosted on the ETH Research Collection platform, guaranteeing long-term preservation. The code to reproduce all steps (training and evaluation of models, patch creation, downloading raw data) is hosted on the project’s public GitHub repository. The Croissant metadata for the dataset is available through HuggingFace.
The DOI of the dataset is 10.3929/ethz-b-000674193.
@@
7.2 Patches creation
In this section, we describe in more details the implementation of the data assembly and creation of the dataset. We refer to this step as patch creation, referring to its final product, square image patches.
As evidenced in Table 3, the various data sources are not aligned in terms of resolution. The first step in the patches creation process is to upsample the data sources to 101010\,10m resolution when necessary. For the continuous variables (the 202020\,20m and 606060\,60m Sentinel-2 bands, the ALOS-2 PALSAR-2 bands, the ALOS DEM band, and the LC probability band) we apply bilinear interpolation. For categorical variables (the Sentinel-2 SCL band, and the LC band) we apply nearest neighbour.
Table 3: Data sources
GEDI S2 ALOS-2 PALSAR-2 ALOS DEM LC CH
@@ 7.2 Patches creation
Temporality - monthly yearly once once yearly
The next step is to actually create the patches. We iterate (in parallel) over the Sentinel-2 tiles covering the regions of interest and load the GEDI footprints that fall within the tile. We then iterate over the footprints, and for each footprint, we consider all Sentinel-2 products for the tile at hand in the six months preceding the footprint acquisition, and take the closest one. For the yearly available data products, we take the tile-cropped data from the year of the footprint acquisition. For the fixed data products (DEM and LC) we simply take the tile-cropped map. For all data sources, we crop the data to a $25\times 25$ pixel square centered around the footprint, which we store in a HDF5 file (the dataset hierarchy can be found in Figure 7). At extraction time, the window size of the patch is fixed ($25\times 25$) but users of the dataset can load any patch size, from pixel-wise ($1\times 1$) to the original $25\times 25$, as long as the patch size is of size $12k+1\times 2k+$ 1 to respect the central pixel approach.
We then define a train/validation/test split. We do not perform the split at the GEDI footprint level, but at the Sentinel-2 tile level, meaning that all patches within the same tile will belong to the same split. To align with our geographical distribution and vegetation types distribution, we define a train/validation/test per region and enforce that the vegetation types distribution within each split should match the whole region’s distribution. This is done through a random exploration within the space of possible combinations, where a combination is selected if its distribution is deemed sufficiently close to the original distribution; and where in the end we select the closest combination out of all the selected ones. We assign $65\%$ of the tiles for training, $15\%$ for validation, and the remaining $20\%$ for testing. To which split each tile belongs can be found on our project’s website, you can also download the mapping directly by running the following command:
wget "https://libdrive.ethz.ch/index.php/s/VPio6i5UlXTgir0/download?path=%2F &files=biomes_splits_to_name.pkl&downloadStartSecret=gxairgqzc" -O biomes_sp lits_to_name.pkl
@@ 7.3 GBDT model initialization parameters
7.4 Dataset hierarchy
In Figure 7, we provide an example of how the data is structured within an .h5 file. We briefly describe each data field.
Note that in the remainder of this section, we assume that f is as follows: with h5py.File(’data_subset-2019-v4_5-20.h5’,’r’) as f:
  • S2_bands: the Sentinel-2 L2A bands (except SCL), in their pre-processed digital number format. The order of the bands (B01, B02, B03, B04, B05, B06, B07, B08, B8A, B10, B11, B12) is stored in the attributes, and can be accessed vif[’21JXM’][’S2_bands’].attrs[’order’]
  • S2_bands: the Sentinel-2 L2A Scene Classification (SCL) band. It is stored separately from the other bands for optimal memory management, as it has a different data type
  • S2_vegetation_score: we assign a "vegetation score" to each patch, a number between 00 and $100$ that represents the percentage of pixels in the patch that are vegetated
  • S2_date: the encoded date of acquisition of the Sentinel-2 product from which the patch is extracted. The encoding is the number of days between the acquisition date and the start of the GEDI mission (date := acquisition_date - start_of_mission)
  • S2_pbn: the Sentinel-2 product’s Processing Baseline Number (PBN)
  • S2_ron: the Sentinel-2 product’s Relative Orbit Number (RON)
  • S2_boa_offset: binary variable equal to $1$ if the product was acquired after January $25$th, $2022$, and equal to 00 otherwise. This will be used by the data loader to calculate the surface reflectance values from the digital number values
  • agbd: the reference GEDI biomass value
  • agbd: the standard error of the reference GEDI biomass value
  • date: the encoded date of acquisition of the GEDI footprint. Again, the encoding is the number of days between the acquisition date and the start of the GEDI mission
  • lat_offset, lat_decimal, lon_offset, lon_decimal: to store the WGS84 latitude and longitude information more efficiently, we split each coordinate into its fractional (offset) and integral (decimal) parts. We store the fractional part as a single precision floating point number, and the integral part as an unsigned 8-bit integer.
  • ...: the additional GEDI features can be accessed via f[’21JXM’][’GEDI’].keys(), and a description of the features can be found in the GEDI documentation
  • ALOS_bands: the ALOS-2 PALSAR-2 bands, in their pre-processed digital number format. The order of the bands (HH, HV) is stored in the attributes, and can be accessed via f[’21JXM’][’ALOS_bands’].attrs[’order’]
  • LC: the Copernicus Land-Cover bands. The first band holds the labels, while the second one holds the corresponding class probabilities
  • DEM: the Digital Elevation Model
  • ch: the Canopy Height by Lang et al.
  • ch_std: the estimated standard deviation of the Canopy Height
[…]
Figure 7: Sample dataset hierarchy.
7.5 Feature importance plots
The lightgbm framework provides a built-in capability to quantify the importance of each of the input features a model has access to. For each feature, it measures the total gains of splits which use the feature. We compute the feature importance for the GDBT model with all features as input, and report the corresponding plot in Figure 8.
Figure 8: Feature importance for the GBDT model using all input features.
7.6 Confusion density plots
We further provide the confusion density plots for the best performing model of each architecture, in Figure 9. It shows good agreement between the predictions and GEDI reference, and provides insights on the distribution of the biomass values.

### v2

@@ 4.2 Results
All quantitative results are reported in Table 2. The lowest error is achieved by the Lang et al. architecture, trained on all considered input channels but CH. The best models achieve RMSE values slightly below $60$ Mg/ha, consistent with previous literature. We elaborate on various aspects in the following paragraphs.
@@
Role of input features. In order to study the impact of each data source, we conduct a series of experiments. We vary the input features available to each model: all features, only the canopy height (CH), only the 101010\,10m RGB (B02, B03, B04) and NIR (B08) S2 bands (RGBN), all S2 bands (S2 (all)), all S2 bands plus the LC map (LC), all S2 bands plus the DEM (DEM), all S2 bands and the two ALOS-2 PALSAR-2 bands (P2), and all features except for the CH. Per-pixel latitude and longitude are always provided as inputs. In each setting, the $5$ models with the highest validation scores are applied to the held-out test set to obtain the mean RMSE values and associated standard deviations reported in Table 2. The best model is the one which has access to additional input features (i.e., on top of S2); we attribute this to the comprehensive, complementary evidence that the model can rely on (namely optical, SAR, elevation, and vegetation).
Patch size. We consider patches of size consistent with the literature [Lang2023, Tolan2024]. Namely, to study how varying the patch-size affects model performance, we train and evaluate with two different patch sizes: $15\times 15$ and $25\times 25$ pixels (see Table 2). Our experiments showed minimal performance differences, favoring the $25\times 25$ pixels patch size. Given the limited spatial context of most existing biomass retrieval methods, it is interesting that increasing the context window yields better results. At present it is unclear whether this is due to real context information or simply a data augmentation effect.
Confusion density plots. We provide the confusion density plots for the best performing model of each architecture, in Figure 5. It shows good agreement between the predictions and GEDI reference, and provides insights on the distribution of the biomass values.
@@
Binned residuals analysis. Expanding on the previous analysis, we also analyse the residual distribution across various AGB intervals, see Figure 6. Besides our baseline models described in Section 4.1 we also include the ESA CCI AGB map, upsampled to 101010\,10m resolution, in this analysis. That comparison is directly relevant, since ESA CCI also uses GEDI to calibrate their model, according to their Algorithm Theoretical Basis Document [ESACCI, p.23]. We point out that the distribution of AGB values is skewed: most ($50$%) of all GEDI labels fall into the first bin ($[0,50]$).
Figure 6: Binned test residuals for the best-performing model of each architecture, and for the ESA CCI predictions.
Per biome residuals analysis. To emphasize how much AGB estimates depend on the biome at hand, we plot the residuals across the biomes, in Figure 7.
Figure 7: Per-biome test residuals for the best-performing model of each architecture, and for the ESA CCI predictions.
Saturation analysis. A known issue for AGB estimation is saturation, which happens when beyond a certain AGB threshold, the RS information no longer reflects the change in AGB, making it challenging to estimate high values. To investigate whether including additional data sources mitigates with this issue, we include the binned residuals of the Lang et al. models trained on the various feature combinations. The results are shown in Figure 8, and indicate that having access to the additional features makes the AGB estimates better particularly in the higher bins.
Figure 8: Binned mean test RMSE ($\downarrow$) and associated standard deviation per model, with various inputs.
@@
× 69.88 $\pm$0.17 69.50 $\pm$0.13 69.49 $\pm$0.11 69.05 $\pm$0.16 68.63 $\pm$0.15 68.43 $\pm$0.10 68.04 $\pm$0.09
× 74.74 $\pm$0.23 65.35 $\pm$0.16 65.06 $\pm$0.26 63.12 $\pm$0.30 62.30 $\pm$0.46 61.73 $\pm$0.30 60.39 $\pm$0.22
× 70.93 $\pm$0.35 64.16 $\pm$0.22 63.92 $\pm$0.40 61.87 $\pm$0.14 60.84 $\pm$0.10 60.46 $\pm$0.15 60.68 $\pm$0.46
× × 67.52 $\pm$0.15 62.57 $\pm$0.22 62.48 $\pm$0.28 61.17 $\pm$0.23 60.28 $\pm$0.42 59.72 $\pm$0.28 58.80 $\pm$0.08
× × 68.38 $\pm$0.15 63.30 $\pm$0.24 63.92 $\pm$0.34 61.64 $\pm$0.12 60.71 $\pm$0.21 60.05 $\pm$0.13 59.03 $\pm$0.09
× × 68.69 $\pm$0.27 62.19 $\pm$0.35 61.94 $\pm$0.21 60.40 $\pm$0.14 59.20 $\pm$0.18 59.08 $\pm$0.05 58.04 $\pm$0.13
× × × × 65.63 $\pm$0.08 61.47 $\pm$0.21 61.34 $\pm$0.03 59.85 $\pm$0.12 59.01 $\pm$0.05 58.64 $\pm$0.08 57.87 $\pm$0.11
× × × × × 63.64 $\pm$0.06 61.29 $\pm$0.13 61.22 $\pm$0.27 60.02 $\pm$0.18 59.47 $\pm$0.26 58.67 $\pm$0.08 58.16 $\pm$0.19
Table 2: Mean test RMSE ($\downarrow$) and associated standard deviation per model, with various inputs. Crosses denote the presence of a feature, values in brackets denote patch size. Values are colored from dark red (higher) to dark blue (lower).
Figure 9: Sentinel-2 tile 30NXM (Ghana), our best model’s prediction, the corresponding ESA CCI map, and GEDI L4B product. Global view (top) and zoomed in views (bottom).
Qualitative results inspection. For a visual impression, Figure 9 shows the predictions of the best-performing model versus the CCI map for a selected Sentinel-2 tile (30NXM, located in Ghana). As the GEDI L4A reference footprints are too small and sparse to visualize, we include the GEDI L4B 1km product [GEDI_L4B], which provides statistically inferred estimates of mean AGB from the footprints located within the borders of each 1 km cell. This figure illustrates how the high spatial resolution of our estimates is able to preserve spatial details that are not captured in the CCI product, meaning that they would not be restored by conventional upscaling of existing maps.
@@ 5 Discussion
Results. Our various analyses indicate that including input features beyond Sentinel-2 is beneficial for the task of AGB estimation, and that models with more parameters tend to better capture the complex relationship between input features and AGB.
Stand-alone task. As AGB is highly dependent on CH (via allometric equations), one could wonder whether CH information is enough to estimate AGB. We conclude that it is not the case: our results show that learning AGB "from scratch" (using S2, P2, LC and DEM) yields better results than learning from CH directly. While this applies to the specific Lang et al. CH product at hand, the most recent ones published [Pauls2024, Tolan2024] also rely on GEDI CH estimates, so we expect similar results. We thus argue that AGB estimation should be tackled as its own task.
Limitations. Despite the nominal 10 m resolution of our estimates, the effective spatial resolution at which biomass is identified is lower. Additionally, the uncertainty in the footprint-level GEDI AGBD product used as reference data influences the accuracy and uncertainties of the model predictions. We elaborate on these points in the Appendix, Section 8.5.
@@ 6 Conclusion
We provide a ML-ready, unrestricted, easily accessible dataset for high-resolution biomass estimation from remote sensing data. This dataset consists of $\approx 16\cdot 10^{6}$ patches, covering a diverse and representative selection of geographic locations with a total of $\approx 500\cdot 10^{6}$ hectares and two different years ($2019$ and $2020$). Furthermore, we complement the dataset with a suite of baseline architectures and trained weights for others to build upon. We hope that our dataset may help to reduce the geographical and ecosystem-related biases of biomass models, enhance the accuracy of future biomass maps, and perhaps contribute towards a regular, operational high-resolution monitoring system.
@@ 7 Acknowledgments
8 Appendix
8.1 Dataset description
Each sample in the dataset contains a pair of pre-cropped, pre-normalized images along with their corresponding AGB labels.
The load_dataset function provides the following configuration options:
  • normalize_data: {True, False}
    Whether to return normalized (0-1) data or raw values.
  • additional_features: {}
    A list of additional features the dataset should include. Refer to the documentation below for more details.
  • patch_size: 25
    The size of the returned patch (in pixels). The maximum value is 25 pixels, which corresponds to 250 meters.
Each image consists of 24 channels, organized into the following categories:
  • Spectral Bands:
    B01, B02, B03, B04, B05, B06, B07, B08, B8A, B09, B11, B12
  • Geographical Coordinates:
    lat_cos, lat_sin, lon_cos, lon_sin, where:
    lat_cos $=(cos(\pi*lat/90)+1)/2$
    lat_sin $=(sin(\pi*lat/90)+1)/2$
    lon_cos $=(cos(\pi*lon/180)+1)/2$
    lon_sin $=(sin(\pi*lon/180)+1)/2$
  • ALOS PALSAR Bands:
    alos_hh, alos_hv
  • Canopy Heights:
    ch, ch_std
  • Land Cover Information:
    lc_cos, lc_sin, lc_prob, where:
    lc_cos $=(cos(2\pi*lc\_class/100)+1)/2$
    lc_sin $=(sin(2\pi*lc\_class/100)+1)/2$
    lc_prob $=lc\_prob/100$
  • Digital Elevation Model (DEM):
    dem
The channels are ordered as follows: (Spectral Bands) | (Geographical Coordinates) | (ALOS PALSAR Bands) | (Canopy Heights) | (Land Cover Information) | DEM. In more details, this corresponds to the following bands: (B01 B02 B03 B04 B05 B06 B07 B08 B8A B09 B11 B12) | (lat_cos, lat_sin, lon_cos, lon_sin) | (alos_hh, alos_hv) | (ch, ch_std) | (lc_cos, lc_sin, lc_prob) | dem.
Users can include a list of additional features from the options below in their dataset configuration:
  • "agbd_se" - AGBD Standard Error: The uncertainty estimate associated with the aboveground biomass density prediction for each GEDI footprint.
  • "elev_lowes" - Elevation: The height above sea level at the location of the GEDI footprint.
  • "leaf_off_f" - Leaf-Off Flag: Indicates whether the measurement was taken during the leaf-off season, which can impact canopy structure data.
  • "pft_class" - Plant Functional Type (PFT) Class: Categorization of the vegetation type (e.g., deciduous broadleaf, evergreen needleleaf).
  • "region_cla" - Region Class: The geographical area where the footprint is located (e.g., North America, South Asia).
  • "rh98" - RH98 (Relative Height at 98%): The height at which 98% of the returned laser energy is reflected, a key measure of canopy height.
  • "sensitivity" - Sensitivity: The proportion of laser pulse energy reflected back to the sensor, providing insight into vegetation density and structure.
  • "solar_elev" - Solar Elevation: The angle of the sun above the horizon at the time of measurement, which can affect data quality.
  • "urban_prop" - Urban Proportion: The percentage of the footprint area that is urbanized, helping to filter or adjust biomass estimates in mixed landscapes.
  • "gedi_num_days" - Date of GEDI Footprints: The specific date on which each GEDI footprint was captured, adding temporal context to the measurements.
  • "s2_num_days" - Date of Sentinel-2 Image: The specific date on which each Sentinel-2 image was captured, ensuring temporal alignment with GEDI data.
  • "lat" - Latitude: Latitude of the central pixel.
  • "lon" - Longitude: Longitude of the central pixel.
@@ 8.2 Hosting, licensing, and maintenance plan
The raw data, dense predictions, and model weights are hosted on the ETH Research Collection, guaranteeing long-term preservation.
@@
GEDI Sentinel-2 (S2) ALOS-2 PALSAR-2 (P2) ALOS Digital Elevation Model (DEM) Land Classification (LC) CH
@@
Temporality - $\approx$ every 5 days yearly once yearly yearly
Table 3: Data sources
@@ 8.3 GBDT model initialization parameters
8.4 GEDI AGB Estimates
The GEDI instrument uses a LiDAR laser to repeatedly scan a 25m surface footprint, where each measurement consists of a waveform from the laser’s returned signal. From the cumulative return energy, the GEDI team identifies the ground return, which is used as the reference for the relative height (RH) metrics. The RH[X] metrics correspond to the relative height at which [X] percent of the total accumulated energy is returned. From the RH metrics, the GEDI team estimates biomass. To do so, they developed their own allometric equations, based on the Plant Functional Type (PFT), the region of the world, and the RH metrics. In order to calibrate them, they leveraged world-wide crowd-sourced data. Coincident airborne LiDAR and ground plot field inventory were put together. The airborne LiDAR data was used to simulate GEDI-like waveforms and derived RH metrics. These simulated RH metrics were then used to calibrate the biomass equations with the field biomass values as reference data. Further information can be found in the Algorithm Theoretical Basis Document[Kellner2023].
8.5 GEDI Limitations
Resolution. The GEDI mission operates with a $25$-meter diameter footprint on the ground. The laser pulse from GEDI’s LiDAR system follows a Gaussian intensity profile, with the highest precision at the center of the footprint, gradually decreasing toward the edges [Hancock2019]. Within this $25$-meter footprint, GEDI measures the highest reflecting surface, which is then used to derive Relative Height metrics, from which they estimate the AGB. We then chose to rasterize these footprints at a 10-meter resolution, which is a well-established method as supported by the literature [Pauls2024, Lang2023], and aligns with the highest resolution available among the sensors we considered, such as the $10$-meter bands from Sentinel-2.
Uncertainty. The uncertainty in the footprint-level GEDI AGBD product used as reference data influences the accuracy and uncertainties of the model predictions. Future work could incorporate field measurements or airborne LiDAR data. However, given the sparsity of GEDI measurements, co-registering these with in-situ sparse measurements is challenging. Notably, in order to assess the quality of GEDI AGB estimates, the authors of [Hunka2023] had to rely on an aggregated 1km x 1km grid GEDI product (where the samples present within the borders of each 1 km cell are used to statistically infer mean AGBD). Our denser and higher resolution estimates provide a crucial step forward, as well-calibrated estimates can be used as a proxy to assess the uncertainties in GEDI data by comparing them in-situ measurements. We believe that this will ultimately need a long-term community initiative across ecology, forestry, remote sensing and ML.
Geolocation error. While the geolocation error was an issue with the first version of the data released by GEDI, they released an updated version of their dataset (that we are using) with corrected geolocation errors. Notably, the authors of [Tang2023] investigated the geolocation error of the updated version, and “did not find any significant relationship between geolocation error and performance”. They find that 80.8% of the footprints have a geolocation error strictly smaller than 10m.
Sampling pattern Another point to keep in mind is that the spatial distribution of GEDI footprints is not uniform: the sampling gets progressively denser as one approaches towards the borders of the observed region at ±$\pm 51.6^{\circ}$.